\begin{thebibliography}{99}


\bibitem{AN}
J.J.R. Aguirre and V.G.L. Neumann,
{\em Existence of primitive $2$-normal elements in finite fields},
Finite Fields Appl. 73 (2021), 101864.
170--183.

\bibitem{carlitz2} L. Carlitz, {\em Distribution of primitive roots in a finite field} , Quart. J. Math. Oxford Ser. (2) 4(1)
(1953), 4--10.

\bibitem{carlitz} L. Carlitz, {\em Primitive roots in a finite field}, Transactions of the American Mathematical Society 73 (1952) 373-382.

%\bibitem{CH} S. D. Cohen and S. Huczynska, {\em The primitive normal basis theorem without a computer}, Journal of London %Mathematical Society, v. 67, n. 1, pp. 41--56, 2003.

\bibitem{davenport1} H. Davenport, {\em Bases for finite fields}, Journal of the London Mathematical Society 43, pp. 21--39.

\bibitem{davenport2} H. Davenport, {\em On primitive roots in finite fields}, Quart. J. Math. Oxford 8(1) (1937), 308--312.

\bibitem{diffie} W. Diffie and M.E. Hellman, {\em New directions in cryptography}, IEEE Trans. Inf. Theory  22,  pp. 644--654, 1976.

\bibitem{gao} S. Gao, {\em Normal basis over finite fields}, PhD thesis, University of Waterloo (Canada), 1993.


\bibitem{knormal} S. Huczynska, G.L. Mullen, D. Panario and D. Thomson,
{\em Existence and
	properties of k-normal elements over finite fields}, Finite Fields Appl. 24 (2013),
170--183.

\bibitem{katz} Nicholas M. Katz, {\em An Estimate for Character Sums},
Journal of the American Mathematical Society, Vol. 2, No. 2. (Apr., 1989), pp. 197-200

\bibitem{lenstra} 
H. W. Lenstra, R. Schoof, Primitive normal bases for finite fields, Mathematics of Computation 48 (1987) 217-231.


\bibitem{Nied} R. Lidl, H. Niederreiter, {\em Finite Fields: Encyclopedia of Mathematics and Its Applications}, vol. 20, 2nd ed., Cambridge University Press, Cambridge, 1997.



\bibitem{lucas} L. Reis:
{\em Existence results on $k$-normal
elements over finite fields},
Rev. Mat. Iberoam. 35(3) (2019),
805--822

\bibitem{lucas1}
L. Reis and D. Thompson, {\em Existence of primitive $1$-normal elements in finite fields}, Finite Fields and Their Applications 51 (2018), 238--269.

\bibitem{rani} M. Rani, A. K. Sharma, S. K. Tiwari and A. Panigrahi, {\em On $r$-primitive $k$-normal elements  with prescribed norm and trace over finite fields}, Finite Fields and Their Applications 91 - 102253, 2023.

%\bibitem{panario} G.L. Mullen, D. Panario (Eds.), Handbook of Finite Fields, CRC Press, Boca Raton, FL, 2013.

\bibitem{SAGE} W. A.  Stein et al., {\em Sage Mathematical Software}, www.sagemath.org, version 8.1, 2017.
\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage[latin1]{inputenc}
\usepackage{indentfirst}
\usepackage{amsfonts}
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{latexsym}
\usepackage{amscd}
\usepackage[all]{xy}
\usepackage{ulem}

\usepackage{hyperref}
\usepackage{xcolor}
\usepackage[linesnumbered,lined,ruled,commentsnumbered]{algorithm2e}

\newcommand{\F}{\mathbb {F}}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{definition}[theorem]{Definition}
\newtheorem{lemma}[theorem]{Lemma}

\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{remark}[theorem]{Remark}

\def \ord {\operatorname{ord}}
\def \Ord {\operatorname{Ord}}

\begin{document}

\title[Existence of primitive $k$-normal elements for critical values over finite fields]{Existence of primitive $k$-normal elements for critical values over finite fields}

\author{Victor G.L. Neumann, Josimar J. R. Aguirre and Sarah F. M. Mazzini}

\maketitle

\vspace{1ex}
\small{Instituto de Matem\'{a}tica e Estat\'{i}stica, Universidade Federal de Uberl\^{a}ndia, 
	Av. J. N. \'{A}vila 2121, 38.408-902 Uberl\^{a}ndia -MG, Brazil }

\vspace{8ex}
\noindent
\textbf{Keywords:} Finite fields, normal elements, $k$-normal elements.\\
\noindent
\textbf{MSC:} 12E20, 11T30.

\begin{abstract}
Let $\mathbb{F}_{q^n}$ be a finite field with $q^n$ elements. An element $\alpha \in \mathbb{F}_{q^n}$ is called $k$-normal over $\mathbb{F}_q$ if $\alpha$ and its conjugates generate a vector subspace of $\mathbb{F}_{q^n}$ of dimension $n-k$ over $\mathbb{F}_q$. The existence of primitive $k$-normal elements and related properties have been studied throughout the past few years for $k > n/2$.
In this paper, we provide general results on the existence of primitive $k$-normal elements for the critical value $k = n/2$, which have not been studied until now, except for $n = 4$.
Furthermore, we show the strength of this result by providing a complete characterization of the existence of primitive $3$-normal elements in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$.
\end{abstract}

\maketitle

\section{Introduction}
Let $\mathbb{F}_{q^n}$ be a finite field with $q^n$ elements, where $q$ 
is a prime power and $n$ is a positive integer.
The field $\mathbb{F}_{q^n}$ has interesting structures related to the two basic field operations.
Regarding the multiplicative structure, $\mathbb{F}_{q^n}^*$ is cyclic and any generator $\alpha \in \mathbb{F}_{q^n}^*$ is called \textit{primitive element}.
On the other hand, in the additive structure, we can consider $\mathbb{F}_{q^n}$ as an $\mathbb{F}_q$-vector space. 
If for an element $\alpha \in \mathbb{F}_{q^n}$ the set $B_{\alpha} = \{ \alpha^{q^i} \mid 0 \leq i \leq n-1 \}$ forms an $\mathbb{F}_q$-basis of $\mathbb{F}_{q^n}$, then $B_\alpha$ is called a normal basis and $\alpha$ is called a \textit{normal element} over $\mathbb{F}_q$.

Normal bases are frequently used in cryptography and computer algebra systems
due to their efficient exponentiation properties
(see \cite{gao} for an overview on normal elements and their applications).
Primitive
elements are constantly used in cryptographic applications, such as in the discrete
logarithm problem (see \cite{diffie}).

By combining these two properties, namely \textit{primitive normal element}, we can study the multiplicative
structure of $\mathbb{F}_{q^n}$ while simultaneously viewing $\mathbb{F}_{q^n}$ as a vector space over $\mathbb{F}_q$. The \textit{Primitive Normal Basis Theorem} states that for any extension field $\mathbb{F}_{q^n}$ of $\mathbb{F}_q$, there exists a basis composed of primitive normal elements. This result was first proved by Lenstra and Schoof \cite{lenstra} using a combination of additive and multiplicative character sums, sieving results, and a computer search.

In $2013$, Huczynska et al. \cite{knormal} introduced the concept of $k$-normal elements as an extension of the usual definition of normal elements.

\begin{definition}
Let $\alpha \in \mathbb{F}_{q^n}$. We say that $\alpha$ is a $k$-normal element of $\mathbb{F}_{q^n}$ over $\mathbb{F}_q$
if the set $B_{\alpha} = \{ \alpha^{q^i} \mid 0 \leq i \leq n-1 \}$
generates a vector subspace of $\mathbb{F}_{q^n}$ of dimension $n-k$ over $\mathbb{F}_q$.
\end{definition}


From the above definition, elements which are normal in the usual sense are $0$-normal 
and from the Primitive Normal Basis Theorem, we know that primitive $0$-normal elements always exist \cite{Nied}.
This definition opened a new line of research on the existence of primitive $k$-normal elements (see \cite{lucas1}, \cite{AN} for results in the cases $k=1$ and $k=2$, respectively).

The general case of the existence of primitive $k$-normal elements is discussed in \cite{lucas}. Although the number of works in this line of research is extensive (see, for example, \cite{rani} and the references therein), in all cases, $k$-normality is only discussed for $k>\frac{n}{2}$ due to an inequality 
requiring $\frac{n}{2} - k$ to be positive.

In \cite{AN}, the authors determine the prime powers $q$ for which $\mathbb{F}_{q^n}$ has
primitive $2$-normal elements over $\mathbb{F}_q$. The special case $n=4$ is treated differently and motivated this work, in which we generalize the ideas used in \cite{AN}  and discuss the case $k=\frac{n}{2}$, obtaining algebraic and analytic conditions for the existence of primitive $k$-normal elements in $\mathbb{F}_{q^{n}}$ over $\mathbb{F}_q$. In particular, we give a complete answer for the case $k=3$.

The conditions for the existence of primitive $k$-normal elements depend on some estimates involving character sums, whose upper bounds are of the order $O\left( q^{n/2} \right)$, where the sums are taken over all elements of the finite field $\mathbb{F}_{q^n}$. These conditions are as follows (see \cite{AN}, \cite{lucas}).


\begin{theorem}\label{teo-exist}
Let $f \in \mathbb{F}_q[x]$ be a divisor of $x^n-1$ of degree $k$. If
$$
q^{n/2-k} \geq W(q^n-1)W(x^n-1),
$$
then there exist primitive $k$-normal elements in $\mathbb{F}_{q^n}$, where $W(t)$ denotes the number 
of square-free (monic) divisors of $t$, for $t$ being either a positive integer or a monic polynomial over $\mathbb{F}_q$.
\end{theorem}

In the above theorem, it is clear that the result cannot be applied when $n=2k$. This case yields what we call \textit{critical values} for $n$ and $k$. Thus, we need a new approach to deal with this case, which will be shown throughout this paper as follows: In Section \ref{preliminares}, we provide some background material. In Section \ref{cv}, we present general conditions for the existence of primitive $k$-normal elements in $\F_{q^n}$ over $\F_q$ for the critical values, generalizing the approach used in \cite{AN}. Finally, in Section \ref{pc}, we apply the results from previous sections to provide a complete characterization of the existence of primitive $3$-normal elements in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$.
In the Appendix, we show the SageMath procedures used in this paper.
%
%
%

\section{Preliminaries}\label{preliminares}

Throughout this paper, $q$ denotes a prime power, $\mathbb{F}_q$ denotes the finite field with $q$ elements, and $\mathbb{F}_{q^n}$ denotes its extension of degree $n$.

In this section, we present definitions and results that will be useful in subsequent sections.

\subsection{Linearized polynomials and the $\mathbb{F}_q$-order}

Here we present some definitions and basic results on linearized polynomials over finite fields that are frequently used in this paper. 

\begin{definition}
Let $f \in \mathbb{F}_q[x]$ a polynomial of the form $f(x)=\sum_{i=0}^r a_ix^i$.
\begin{enumerate}
\item[(a)] The polynomial $L_f(x)=\sum_{i=0}^r a_ix^{q^i}$ is the 
linearized $q$-associate of $f$.
\item[(b)] For $\alpha \in \mathbb{F}_{q^n}$, we set 
$L_f(\alpha)= \sum_{i=0}^r a_i \alpha^{q^i}$.
\end{enumerate}
\end{definition}

The polynomial 
$L_f$ 
can be viewed as a linear transformation $L_f: \mathbb{F}_{q^n} \longrightarrow \mathbb{F}_{q^n}$ over $\F_q$ that also has additional properties.

\begin{lemma}[\cite{Nied}, Lemma 3.59]
Let $f,g \in \mathbb{F}_q[x]$. The following hold:
\begin{enumerate}
\item[(a)] $L_f(x)+L_g(x)=L_{f+g}(x)$.
\item[(b)] $L_{fg}(x)=L_f \left( L_g(x) \right) = L_g \left( L_f(x) \right)$.
\end{enumerate}
\end{lemma}


\begin{definition}\label{Order}
Let $g \in \F_q[x]$ be a monic polynomial.
We say that an element $\alpha \in \F_{q^n}$ has $\F_q$-order $g$
if $g$ is the lowest degree monic polynomial such that $L_g(\alpha)=0$. In this case, we denote it
by $g=\text{Ord}(\alpha)$.
\end{definition}
It is known that the $\F_q$-order of an element $\alpha \in \F_{q^n}$
divides $x^n-1$.
We also have the following equivalences. 

\begin{theorem}[\cite{knormal}, Theorem 3.2]\label{equiv-knormal}
 Let $\alpha \in \mathbb{F}_{q^n}$. The following three properties are equivalent:
\begin{enumerate}
\item[(i)] $\alpha$ is $k$-normal over $\F_q$.
\item[(ii)] If $V_{\alpha}$ is the $\F_q$-vector space generated by $\{ \alpha, \alpha^q, \ldots, \alpha^{q^{n-1}} \}$, then $\dim V_{\alpha}$ is $n-k$.
\item[(iii)] $\alpha$ has $\mathbb{F}_q$-order of degree $n-k$.
\end{enumerate} 
\end{theorem}

\subsection{Freeness and Characters.}

The concept of \textit{freeness} is useful in the construction of certain characteristic functions over finite fields. This concept was introduced in Carlitz \cite{carlitz} and Davenport \cite{davenport1},
and refined in Lenstra and Schoof \cite{lenstra}. 

\begin{definition}
Let $m$ be a divisor of $q^n-1$. An element $\alpha \in \mathbb{F}_{q^n}^*$ is called $m$-free if for every divisor $d$ of $m$ with $d > 1$, $\alpha$ is not a $d$-th power in $\mathbb{F}_{q^n}^*$. 
\end{definition}


It is well known that an element $ \alpha \in \mathbb{F}_{q^n}^*$ is primitive if and only if $\alpha$ is $(q^n-1)$-free.


Consider the multiplicative structure of $\mathbb{F}_{q^n}$. Let $m$ be a divisor of $q^n-1$ and define
\[
\int \limits_{d|m} \eta_d = \sum_{d|m} \frac{\mu(d)}{\varphi(d)} \sum_{(d)} \eta_d,
\]
where $\eta_d$ is a multiplicative character of $\F_{q^n}$, and the sum
$ \sum_{(d)} \eta_d$ runs over all multiplicative characters of order $d$.
It is known that there exist $\varphi(d)$ of these characters.

\begin{proposition}\label{caracteristica-mlivre}
Let $m$ be a divisor of $q^n-1$, and $\theta(m)= \dfrac{\varphi(m)}{m}$.
For any $\alpha \in \F_{q^n}$ we have
 $$w_m(\alpha) = 
\theta(m) \int_{d|m} \eta_d(\alpha)
=\left\{
\begin{array}{ll}
1 \quad & \text{if } \alpha \text{ is } m\text{-free}, \\
\theta(m) & \text{if } \alpha=0 ,\\
0            & \text{otherwise.}
\end{array}
\right.
$$
\end{proposition}
\begin{proof}
See \cite[section 5.2]{knormal} or \cite[Theorem 2.15]{lucas}.
\end{proof}

To finish this section, we present an estimate that is used in the next section.

\begin{lemma}[\cite{katz}, Theorem 1]\label{cotaenFq}
Let $F$ be a finite field,
$n \geq 1$ be an integer and $E$ be an extension field of $F$ of degree $n$.
Let $\chi$ be any nontrivial complex-valued multiplicative character of $E^{\times}$ (extended by zero to all of E ), and $x \in E$ any element that generates $E$
over $F$. Then
$$
\left| \sum \limits_{t \in F} \chi(t-x) \right| \leq (n-1) \sqrt{\# (F)}.
$$

\end{lemma}

\section{Existence conditions for critical values}\label{cv}

Throughout this section, $k$ is a positive integer, $n=2k$ and
$\alpha \in \mathbb{F}_{q^n}$ is a normal element over $\mathbb{F}_{q}$. 

\begin{proposition}\label{prop1}
Let $g \in \mathbb{F}_q[x]$ be a monic divisor of $x^n-1$ with $\deg g=k$, and consider $\beta=L_f(\alpha) \in \mathbb{F}_{q^n}$ where $f=\frac{x^n-1}{g}$.
If $x-1$ divides $g$ then there exists $u_0 \in \mathbb{F}_q$ such that
$\Ord(\beta + u) =g$
for all $u \in \mathbb{F}_q \setminus \{u_0\}$. 
\end{proposition}

\begin{proof}
Note that, from Theorem \ref{equiv-knormal}, $\beta=L_f(\alpha)$ is a $k$-normal element over $\mathbb{F}_q$, and we need to find an element $u_0 \in \mathbb{F}_q$ such that $\beta+u$ is also a $k$-normal element for all $u \in \mathbb{F}_q \setminus \{u_0\}$.

Since $(x-1) \mid g$, thus $L_g(u)=0$ and $L_g(\beta+u)=0$, which implies that $\Ord(\beta+u) \mid g$. Suppose that $g=(x-1)^m \cdot h$, with $(x-1) \nmid h$. We analyze two cases: First, if $m \geq 2$, we have
$L_{(x-1)^{m-1}h}(\beta+u)=L_{(x-1)^{m-1}h}(\beta) \neq 0$, implying $\Ord(\beta+u)=(x-1)^m \cdot h_1$, with $h_1 \mid h$. But if $h_1 \neq h$, the degree of $\Ord(\beta)$ is less than $k$, since $L_{(x-1)^m \cdot h_1}(\beta)=L_{(x-1)^m \cdot h_1}(\beta+u)=0$, which is a contradiction. Therefore, $\text{ Ord}(\beta+u)=g$ and $\beta+u$ is $k$-normal for all $u \in \mathbb{F}_q$.

Now we consider $m=1$. In this case, we have $g=(x-1) \cdot h$, with $(x-1) \nmid h$. Since $h(1) \neq 0$, then
$L_h(\beta+u)=L_h(\beta)+L_h(u) = L_h(\beta)+h(1)u\neq 0$ for all $ \in u \in \mathbb{F}_{q^n} \setminus \{u_0\}$ where
$u_0= -h(1)^{-1} L_h(\beta)$.
Therefore, $(x-1) \mid \Ord (\beta+u)$ and, analogously to the previous case, $\Ord(\beta+u)=g$ for all $u \in \mathbb{F}_{q^n} \setminus \{u_0\}$.
\end{proof}

\begin{remark}\label{Remark1}
Let $g \in \mathbb{F}_q[x]$ be a monic divisor of $x^n-1$ with $\deg g=k$, and 
let $\beta\in \mathbb{F}_{q^n}$ be such that $\Ord(\beta)=g$.
Note that $g\neq x^k-1$ if and only if $\mathbb{F}_{q^n}=\mathbb{F}_q(\beta)$. In fact,
$\beta \in \mathbb{F}_{q^d}$ for some  divisor $d$ of $n$ is equivalent to 
$g \mid x^d-1$. 
Since $\deg g=k$ and $d<n$, the last assertion means $g=x^k-1$.

Hence, to get an element $\beta \in \mathbb{F}_{q^n}$ which generates $\mathbb{F}_{q^n}$,
we need that $\Ord(\beta) \neq x^k-1$. Thus, if the hypotheses of Proposition \ref{prop1} are satisfied and $\Ord(\beta) \neq x^k-1$, we may apply Lemma \ref{cotaenFq} in order to prove the existence of a primitive element of the form $\beta+u$, where $u \in \mathbb{F}_q$.

This follows the ideas of Davenport \cite{davenport2} and Carlitz \cite{carlitz2} about the translate property.
\end{remark}

The last remark motivates the following definition.

\begin{definition}
We say that a polynomial $g \in \mathbb{F}_q[x]$ 
satisfies property $(A)$ if $g$ is a monic 
divisor of $x^{n}-1$
of degree $k$ such that
$g\neq x^k-1$ and $(x-1)\mid g$.
\end{definition}

\begin{theorem}\label{teo-principal}
Let $g \in \mathbb{F}_q[x]$ be a polynomial satisfying property (A), $\beta=L_f(\alpha) \in \mathbb{F}_{q^{n}}$ where $f=\frac{x^{n}-1}{g}$, $m$ a divisor of $q^n-1$ and
$N_{\beta}(m)$ 
the number of elements $u \in \mathbb{F}_q$ such that $\beta+u$ is $m$-free.
If
$q^{1/2} \geq (n-1) W(m),$
then $N_{\beta}(m)>0$. In particular, if $q^{1/2} \geq (n-1) W(q^{n}-1)$ there exists 
a primitive $k$-normal element in $\mathbb{F}_{q^{n}}$ over $\mathbb{F}_q$.
\end{theorem}

\begin{proof}
From Remark \ref{Remark1} it follows that $\mathbb{F}_{q^{n}}=\mathbb{F}_q(\beta)$.  Therefore, from Lemma \ref{cotaenFq}, for any non-trivial multiplicative character $\chi$ over $\mathbb{F}_{q^{n}}$, we have
\begin{equation}\label{cotaparag}
 \left| \sum \limits_{u \in \F_q} \chi(\beta+u) \right| \leq (n-1) \sqrt{q}.
\end{equation}
From Proposition \ref{caracteristica-mlivre}, we obtain
\begin{align*}
N_\beta(m) & =\sum_{u \in \F_q} w_m\left( \beta+u \right) = 
\theta(m) \left( \sum_{ u \in \F_q} \chi_1(\beta+u) + \int \limits_{d|m, \ d \neq 1} \sum_{ u \in \F_q} \chi_d(\beta+u) \right).
\end{align*}
Using inequality \eqref{cotaparag}, we get the estimate
\begin{align*}
\left| \frac{N_\beta(m)}{\theta(m)} - q \right| & \leq \left |  \sum_{\substack{d|m \\ d \neq 1}}  
\dfrac{\mu(d)}{\varphi(d)} \sum_{(d)} \sum_{u \in \F_q} \chi_d(\beta+u) \right| \\
& \leq (n-1) \sqrt{q} \sum_{\substack{d|m \\ d \neq 1}} |\mu(d)|
= (n-1)(W(m)-1)\sqrt{q}.
\end{align*}
Thus, \( \frac{N_\beta(m)}{\theta(m)} \geq q - (n-1)(W(m)-1)\sqrt{q} \), which establishes the desired result.
\end{proof}

\begin{remark}\label{no-existe}
Note that the nonexistence of a polynomial $g$ satisfying property $(A)$ implies that $x^{n}-1$ has only $x^k-1$ and $x^k+1$ as monic factors of degree $k$. Moreover, when $x^{n}-1$ has only these two monic degree-$k$ factors, then $\mathbb{F}_{q^{n}}$ cannot contain primitive $k$-normal elements.
Let us prove this claim.

Let $\beta = L_{x^k\pm 1}(\alpha) =\alpha^{q^k} \pm \alpha$. Then,
$\beta^{q^k} = \alpha^{q^{2k}} \pm \alpha^{q^k}=\alpha \pm \alpha^{q^k} = \mp \beta$.
Consequently, $\beta^{2(q^k-1)}=1$. Hence, the multiplicative order of $\beta$ divides $2(q^k-1)$ and $\beta$ cannot be primitive.
\end{remark}

Let us now determine for which field extensions $\mathbb{F}_{q^{n}}$ over $ \mathbb{F}_q$  there exists a polynomial $g\in\mathbb{F}_q[x]$ satisfying property $(A)$.
We will prove that this existence depends on both $k$ and $q$.

\begin{lemma}\label{propertyA-even}
Let $q=2^m,$ with $m\geq 1$, $k=2^s\cdot t \geq 2$, with $s\geq 0$, $t$ an odd positive integer, and $d=\ord_t(2^m).$ The polynomial $x^{n}-1\in\mathbb{F}_q[x]$ has a factor $g\in\mathbb{F}_q[x]$ satisfying (A) if and only if $t>1$ and one of the following conditions holds:
\begin{itemize}
    \item [(a)] $t$ is prime, $d=t-1$ and $t\leq 2^{s}+1$;
    \item [(b)] $t$ is prime and $d<t-1$;
    \item [(c)] $t$ has two different prime factors;
    \item [(d)] $t=r^{\ell}$ for some prime $r$ and $\ell \geq 1$ such that $d<r^\ell-r^{\ell-1}$;
    \item [(e)] $t=r^{\ell}$ for some prime $r$ and $\ell \geq 1$ such that $d=r^\ell-r^{\ell-1}$ and
    $r\leq 2^s+1$.
\end{itemize}
\end{lemma}

\begin{proof}
If $t=1$,
the polynomial $x^{n}-1=(x-1)^{2^{s+1}}$ has only $(x-1)^{2^{s}}$ as a monic factor of degree $k=2^s$.
Therefore, from Remark \ref{no-existe}, we conclude that there is no primitive $k$-normal element of $\mathbb{F}_{q^{n}}$ over $\mathbb{F}_q$.

Let's suppose that $t$ is prime and $d=t-1$. From
\cite[Theorem 2.47]{Nied}, the cyclotomic polynomial
$Q_t(x) = x^{t-1}+\dots+x+1$ is irreducible. In this case, the monic factors of
$x^{n}-1$ of degree $k$ are of the form
$$
g_{a,b}(x)=(x+1)^{a}Q_t(x)^b,
$$ 
with $0 \le a,b \le 2^{s+1}$ and $a+(t-1)b=2^s\cdot t$.
If $a=b$, then $g_{a,b}=x^k-1$. Thus, the existence of a polynomial $g=g_{a,b}$
satisfying (A) is equivalent to the existence of integers 
$1 \le a \neq b \le 2^{s+1}$ with $a+(t-1)b=2^s\cdot t$. Observe that all the integer
solutions of this equation are of the form
$(a,b) = (2^s + (t-1) \ell, 2^s - \ell)$, for all $\ell \in \mathbb{Z}$.
The existence of a positive integer solution different from
$(a,b)=(2^s,2^s)$ with $1 \le a \neq b \le 2^{s+1}$ is equivalent to
$t \le 2^s+1$, which proves (a).

Suppose now that $t$ is prime and $d<t-1$.
By \cite[Theorem 2.47]{Nied} we have that 
$Q_t(x)=h_1(x)\cdots h_{\frac{t-1}{d}}(x)$, where $h_1(x), \ldots ,h_{\frac{t-1}{d}}(x)$ are  monic irreducible distinct polynomials of degree $d$.
In this case, all the monic factors of $x^{n}-1$ of degree $k$ are of the form
$$
g_{a}(x)=(x+1)^{a_0}h_1(x)^{a_1}\cdots h_{\frac{t-1}{d}}(x)^{a_{\frac{t-1}{d}}},
$$ 
such that
$a_0+d(a_1+a_2+\dots+a_{\frac{t-1}{d}})=2^s\cdot t$. If $a_0 \ge 1$ and
$a_i\neq 2^s$, for some $i=1,\cdots, \frac{t-1}{d}$, 
the polynomial $g_a(x)$ satisfies (A).
Since $\frac{t-1}{d}>1$, we can choose $a_1=2^s+1$, $a_2=2^s-1$ and $a_i=2^s$ for $i\neq 1,2$.
Thus, in this case, there exists a polynomial $g$ satisfying (A).

Suppose now that $t$ has at least two different prime factors. 
Let $p_1$ and $p_2$ be two prime factors of $t$. It is clear that
$p_1$ and $p_2$ divide $(2^m)^{\frac{(p_1-1)(p_2-1)}{2}}-1$, thus
$\ord_{p_1p_2}(2^m)< \varphi(p_1 p_2)= (p_1-1)(p_2-1)$. This implies that
the polynomial $Q_{p_1p_2}(x)$ is reducible. Hence, there exist 
distinct monic irreducible polynomials $h_1(x)$ and $h_2(x)$ dividing $Q_{p_1p_2}(x)$.
In this case, the polynomial
$g(x)=\frac{x^k-1}{h_2(x)}\cdot h_1(x)$ satisfies (A), since
$Q_{p_1p_2}(x)$ divides $ x^k-1$.
That proves (c).




Suppose that $t=r^{\ell}$, where $r$ is a prime number, $\ell \geq 1$ and $d<r^\ell-r^{\ell-1}$.
Then $Q_{r^\ell}(x)$ is reducible and we proceed as in case (c) to prove (d).

Finally, suppose that $t=r^{\ell}$, where $r$ is a prime number, $\ell \geq 1$ and $d=r^\ell-r^{\ell-1}$.
Thus ${\ord}_{r^v}(2^m)=r^{v}-r^{v-1}$, for all $1\leq v \leq \ell$ and
all $Q_{r^v}(x)$ are irreducible.
In this case, a monic factor of degree $k$ of $x^{n}-1$ is of the form 
$$
g_{a}(x)=(x+1)^{a_0}Q_r(x)^{a_1}\cdots Q_{r^{\ell}}(x)^{a_{\ell}},
$$ 
with $a_0+a_1(r-1)+\dots + a_{\ell}(r^{\ell}-r^{\ell-1})=2^sr^{\ell}$. Define $b_j=a_j-2^s$. Observe that $g_a(x)$ satisfies condition (A) if and only if there exists a nonzero $(\ell+1)$-tuple of integers $(b_0,\dots,b_{\ell})$ satisfying
$|b_i|\leq 2^s$, for $0\leq i\leq \ell$, $b_0\neq -2^s$ and $$b_0+b_1(r-1)+\dots+b_{\ell}(r^{\ell}-r^{\ell-1})=0.$$
Let $(b_0,\dots,b_{\ell})$ be such an $(\ell+1)$-tuple and
$j=\max\{\,i\, |\, b_i\neq 0\}$. Then the last equation becomes 
$$b_0+b_1(r-1)+\dots+b_{j}(r^{j}-r^{j-1})=0.$$
If $r>2^{s}+1$, then 
\begin{eqnarray*}
|b_0+b_1(r-1)+\dots+b_{j-1}(r^{j-1}-r^{j-2})| &\leq & 2^s(1+(r-1)+\dots+(r^{j-1}-r^{j-2}))\\
& =& 2^sr^{j-1}<(r-1)r^{j-1}
\end{eqnarray*}
and $|(r^j-r^{j-1})b_j|\geq r^j-r^{j-1}$. This implies that there is no such $(\ell+1)$-tuple. Thus, no polynomial $g_a(x)$ satisfies (A).

If $r\leq 2^s+1$, we may choose $b_0=r-1\geq1, b_1=-1$ and $b_i=0$ for all $2\leq i\leq \ell$. This choice yields a polynomial $g_a(x)$ satisfying (A), which completes the proof for case (e).
\end{proof}

\begin{lemma}\label{propertyA-odd}
Let $p$ be an odd prime number, $q=p^m$ a prime power of $p$ and $k=2^s\cdot t$ a positive integer, with $s\geq 0$, $t$ an odd positive integer and $k\geq 2.$ The polynomial $x^{n}-1$ has a factor $g\in\mathbb{F}_q[x]$ satisfying (A) if and only if
one of the following conditions holds:
\begin{itemize}
    \item [(i)] $t>1$;
    \item [(ii)] $t=1$ and $m$ is even;
    \item [(iii)] $t=1$, $m$ is odd and $\ord_{2^{s+1}}(p)<2^s$.
\end{itemize}
\end{lemma}
\begin{proof}
If $t>1$, we may choose
$g=(x^{2^s}-1)(x^{2^s (t-1)} - x^{2^s  (t-2)} + \cdots - x^{2^s} + 1)$, which satisfies (A).

Let us now suppose that $t=1$. Consider the factorization of 
$x^{n}-1= x^{2^{s+1}}-1$
into cyclotomic polynomials $$x^{2^{s+1}}-1=\prod_{i=0}^{s+1} Q_{2^i}(x).$$
Note that if $Q_{2^{s+1}}(x)$ is irreducible, then
all cyclotomic polynomials $Q_{2^i}(x)$ for $0\le i \le s$ are also irreducible, and the only monic
factors of $x^{n}-1$ of degree $k$ are
\[
x^k-1 = \prod_{i=0}^{s} Q_{2^i}(x) \quad
\textrm{and}
\quad
x^k + 1 = Q_{2^{s+1}}(x).
\]
Thus, the existence of a polynomial $g$ satisfying $(A)$ depends on the 
factorization of $Q_{2^{s+1}}(x)$.
By  \cite[Theorem 2.47]{Nied} we have that the cyclotomic polynomial $Q_{2^{s+1}}(x)$ is factored
into $\frac{2^s}{\ord_{2^{s+1}}(q)}$ distinct monic irreducible polynomials in $\mathbb{F}_q[x]$ of the same degree 	
$\ord_{2^{s+1}}(q)$. Hence, the polynomial $g$ satisfying $(A)$ exists if and only if $\ord_{2^{s+1}}(q) < 2^s$. 
Using the fact that 
$\ord_{2^{s+1}}(q) = \frac{\ord_{2^{s+1}}(p)}{\gcd \left( m, \ \ord_{2^{s+1}}(p) \right)}$ we obtain the required result.
\end{proof}

\section{Particular case: $n=6$ and $k=3$}\label{pc}

The Primitive 1-Normal Theorem in \cite{lucas1} employed the method of Gauss sums and properties of the trace function to prove the impossibility of finding primitive 1-normal elements in $\mathbb{F}_{q^2}$ over $\mathbb{F}_q$, which corresponds to the critical value $k=1$. Subsequently, the Primitive 2-Normal Theorem in \cite{AN} used 
the bound given in Theorem \ref{teo-exist} and the factorization of $x^n-1$ in conjunction with sieve methods
to prove the existence of primitive 2-normal elements for 
 $n > 4$. The case $n=4$ is handled with similar ideas to those in the previous section.

From Lemmas \ref{propertyA-even} and \ref{propertyA-odd}, 
for $n=6$ and $k=3$,
a polynomial $g\in \mathbb{F}_q[x]$ satisfying (A) exists if and only if $q$ is an odd prime power or a power of $2$ with an even exponent.
Furthermore, Theorem \ref{teo-principal} implies that a primitive $3$-normal element in
$\mathbb{F}_{q^6}$ over $\mathbb{F}_q$ exists if $q$ also satisfies $q^{1/2} \geq 5 W(q^6-1)$.
In order to improve the bounds obtained from this last condition, 
we apply a sieve method to get lower bounds for $q$. The proof of the next result is similar to that of
\cite[Propositions 3, 5]{AN}, and is therefore omitted.

\begin{proposition}\label{crivo}
Let $q$ be an odd prime power or a power of $2$ with an even exponent,
and let $m$ be a positive divisor of $q^6-1$. Let $p_1,\ldots , p_r$ be prime numbers such that
$\operatorname{rad}(q^6-1)=\operatorname{rad}(m)\cdot p_1 \cdots p_r$. Suppose that
$\delta = 1 - \sum_{i=1}^r \frac{1}{p_i}>0$ and let
$\Delta = \frac{r-1}{\delta}+2$. 
If 
$q^{1/2} \ge 5 W(m) \Delta$, then there exists 
a primitive $3$-normal element in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$.
\end{proposition}

Using all the results obtained in this work, we determine for which prime powers $q$  there exist primitive $3$-normal elements in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$.

\begin{theorem}\label{principaln6}
Let $q$ be a prime power. There exists a primitive $3$-normal element in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$ if and only if
$q$ is an odd prime power or a power of $2$ with an even exponent.
\end{theorem}
\begin{proof}
From Remark \ref{no-existe}, if $q$ is a power of $2$ with an odd exponent, then no $3$-normal element
of $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$ is primitive. Thus, from now on, assume that 
$q$ is an odd prime power or a power of $2$ with an even exponent. From Lemmas \ref{propertyA-even} and \ref{propertyA-odd},
in this case,
there exists a polynomial $g\in \mathbb{F}_q[x]$ that satisfies (A).

Let $t$ and $u$ be positive real numbers such that $t+u > 12$, and let
$p_1,\ldots , p_r$ be all prime factors of $q^6-1$ between $2^t$ and $2^{t+u}$.
Choose $t$ and $u$ such that $S_{t,u}<1$, where $S_{t,u}$ is the sum of the inverses of
all prime numbers between $2^t$ and $2^{t+u}$. Let also $r(t,u)$ be the number
of all prime numbers between $2^t$ and $2^{t+u}$. For the primes $p_1, \ldots , p_r$, let
$\delta$, $m$ and $\Delta$ be the numbers given in Proposition \ref{crivo}. Observe that
$r\le r(t,u)$, $\delta \ge 1 - S_{t,u}$ and $\Delta \le 2 + \frac{r(t,u)-1}{1-S_{t,u}}=: \Delta_{t,u}$. From \cite[Lemma 2.9]{AN},
we have $W(m) \le A_{t,u} q^{6/(t+u)}$, where
\[
A_{t,u} = \prod_{\substack{\mathfrak{p}< 2^t \\ \mathfrak{p} \textrm{ is prime}}} \frac{2}{\sqrt[t+u]{\mathfrak{p}}}.
\]
Thus, if $q \ge \left( 5\cdot A_{t,u} \cdot \Delta_{t,u}\right)^{\frac{2(t+u)}{t+u-12}}$, then
\[
q^{1/2} \ge 5 \cdot A_{t,u}\cdot q^{\frac{6}{t+u}} \cdot \Delta_{t,u} \ge
5 W(m) \Delta.
\]
From Proposition \ref{crivo}, this implies the existence of a primitive $3$-normal element in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$.

Taking $t=6.48$ and $u=11.23$, and using SageMath \cite{SAGE}, we get
$r(t,u)=19137$,
$S_{t,u}\cong 0.976788370762892 <1$ and
\[
\left( 5\cdot A_{t,u} \cdot \Delta_{t,u}\right)^{\frac{2(t+u)}{t+u-12}} < 6.46 \cdot 10^{73}.
\]

Now consider $q<M=6.46 \cdot 10^{73}$. Let $u$ denote the number of distinct prime divisors of
$q^6-1$. Since $q^6-1 < M^6-1$, it follows that $u \le 177$.
For any non-negative integer $r \le u$, let $m$ be the product of the first $u-r$ prime factors of
$q^6-1$ and let $\{p_1,\ldots , p_r\}$ be the other prime factors of $q^6-1$.
Let $S_{u,r}$ be the sum of the reciprocals of the first $r$ prime numbers starting from the
$(u-r+1)$th prime, and 
define $\Delta_{u,r}= \frac{r-1}{1-S_{u,r}}+2$.
For the primes $p_1, \ldots , p_r$, let
$\delta$, $m$ and $\Delta$ be the numbers given in Proposition \ref{crivo}.
If $S_{u,r}<1$,
we get
\begin{equation}\label{cota}
5 W(m) \Delta \le 5 \cdot 2^{u-r} \cdot \Delta_{u,r}.
\end{equation}
Let us define
\[
N = \max_{1 \le u \le 177} \{\min_{0 \le r \le u} \{(5 \cdot 2^{u-r} \cdot \Delta_{u,r} )^2\}\}.
\]
From Proposition \ref{crivo} and Inequality \eqref{cota}, we get that
if $q \ge N$ there exists a primitive $3$-normal element in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$.
Using SageMath, from $M=6.46 \cdot 10^{73}$ we obtain $N<1.58 \cdot 10^{11}$.

If we repeat this process several times we get the following. From $M= 1.58 \cdot 10^{11}$, we get
$u \le 39$ and $N<1.10 \cdot 10^{8}$.
From $M= 1.10 \cdot 10^{8}$, we get
$u \le 30$ and $N<3.31 \cdot 10^{7}$.
Finally, from $M= 3.31 \cdot 10^{7}$, we get
$u \le 29$ and $N<2.79 \cdot 10^{7}$.

There are $1,736,412$ prime powers less than $2.79 \cdot 10^{7}$ that are either odd prime powers or powers of $2$ with an even exponent.
Using SageMath, we apply Algorithm \ref{sieve} to these prime powers to verify whether Proposition \ref{crivo} can be used to ensure the existence of primitive $3$-normal elements in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$.
We conclude that for $9,551$ of these prime powers,
the inequality $q^{1/2} \ge 5 W(m) \Delta$ fails for
all factorization of the form $\operatorname{rad}(q^6-1)=\operatorname{rad}(m)\cdot p_1 \cdots p_r$.

To prove the existence of a primitive $3$-normal element in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$ for these $9,551$ prime powers, we apply Algorithm \ref{explicit}, which is based on Theorem \ref{teo-principal}. For these cases, Algorithm \ref{explicit}
finds a primitive $3$-normal element in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$
of the form $L_f(\alpha)+\delta$ except for $q=7$, where $\delta \in \mathbb{F}_q$,
$\alpha$ is a normal element in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$,
and the polynomial $f$ is chosen as follows.

Let $a$ be a primitive element of $\mathbb{F}_{q^6}$, and let $u = a^{(q^6-1)/(q-1)}$ be a primitive element of $\mathbb{F}_q$. If $q = 2^{2m}$ or $q \equiv 1 \pmod{6}$, then $3 \mid (q-1)$, and we set $b = u^{(q-1)/3}$.
In both cases, we have the factorization:
\[
x^6-1 = (x-1)(x+1)(x-b)(x+b+1)(x+b)(x-b-1)
\]
and we choose $f = (x-b)(x+b+1)(x+b) = x^3 + (b+1)x^2 + (b+1)x + b$. Then $g = (x^6-1)/f$ is a polynomial that satisfies (A).
If $q$ is a power of $3$ or $q \equiv 5 \pmod{6}$, then we choose $f = (x+1)(x^2+x+1) = x^3 + 2x^2 + 2x + 1$, and $g = (x^6-1)/f$ is a polynomial that satisfies (A).

For $q=7$, we have $x^6-1=(x - 1) (x + 1)  (x - 2)  (x + 2)  (x - 3)  (x + 3)$
and
any root $\alpha$ of $x^6 + 2x^5 + 2x^4 + 3x^3 + 4x^2 + 2x + 2$ is normal
in $\mathbb{F}_{7^6}$ over $\mathbb{F}_7$. If
$f=(x - 2)(x - 3)(x - 4)$ and $\beta = L_f(\alpha)$, we get that 
$\beta+1$ is a primitive $3$-normal element in $\mathbb{F}_{7^6}$ over $\mathbb{F}_7$.
\end{proof}

\begin{remark}
From the proof of Theorem \ref{principaln6}, it is guaranteed that if $\alpha$ is normal, $f$
is a monic polynomial dividing $x^6-1$, $g = (x^6-1)/f$ satisfies  (A) in $\mathbb{F}_q[x]$, and $q \geq 2.79 \cdot 10^{7}$, then there exists an element $L_f(\alpha) + u$ that is primitive $3$-normal in $\mathbb{F}_{q^6}$ over $\mathbb{F}_q$ for some $u \in \mathbb{F}_q$. Now, also from Theorem \ref{principaln6}, if $q < 2.79 \cdot 10^{7}$, this does not necessarily hold for all monic polynomials $f \mid (x^6-1)$, but there is at least one such polynomial $f$ for which we obtain a primitive $3$-normal element.
\end{remark}

\section*{Appendix: Algorithms in SageMath} \label{apendice}
{\scriptsize
\begin{algorithm}
\KwIn{A prime power $q$}
\KwOut{True or False}
LPrime $\gets$ List of the prime divisors of $q^6 - 1$ in increasing order \\
$u \gets$ Length(Lprime)\\
Value $\gets$ False\\
$r \gets 0$\\
\While{$r \le u$ and Value $=$ False}
    {
    $S\gets$ List of the last $r$ elements of Lprime\\
    $\delta \gets 1- \displaystyle\sum_{p\in S} \frac{1}{p}$\\
    \If{$\delta >0$}
        {
        $\Delta \gets \frac{r-1}{\delta}+2$\\
        $N \gets (5\cdot 2^{u-r} \cdot \Delta)^2$\\
        \If{$q\ge N$}
            {
            Value $\gets$ True
            }
        }
    $r\gets r+1$
    }
\Return Value
\caption{Sieve($q$)}\label{sieve}
\end{algorithm}
}

{\scriptsize
\begin{algorithm}
\KwIn{A prime power $q \neq 2^{2m+1}$}
\KwOut{A primitive $3$-normal element of the extension $\mathbb{F}_{q^6}\mid \mathbb{F}_q$ or False}
$a \gets$ A primitive element of $\mathbb{F}_{q^6}$\\
$\alpha \gets$ A normal element of $\mathbb{F}_{q^6}$ over $\mathbb{F}_{q}$\\
$\gamma \gets$ False\\
\uIf{$q=2^{2m}$ or $q \equiv 1 \pmod 6$}
     {
     $b \gets a^{\frac{q^6-1}{3}}$\\
     $f \gets x^3 + (b+1)x^2 + (b+1)x + b$\\
     $\beta \gets L_f(\alpha)$
     }
\Else{
     $f \gets x^3 + 2x^2 + 2x + 1$\\
     $\beta \gets L_f(\alpha)$
     }
NotEnd $\gets$ True \\
\uIf{$\ord(\beta)=q^6-1$}
    {
     NotEnd $\gets$ False \\
     $\gamma \gets \beta$
    }
\Else{
    $u \gets a^{\frac{q^6-1}{q-1}}$\\
    $j \gets 0$\\
    \While{$j<q-1$ and NotEnd}
          {\If{$\ord(\beta+u^j)=q^6-1$}
               {
               $p \gets \displaystyle\sum_{i=0}^{5} (\beta+u^j)^{q^i}x^{5-i}$\\
               $d \gets \deg (\gcd(p,x^6-1))$\\
               \If{d=3}
                  {NotEnd $\gets$ False \\
                   $\gamma \gets \beta+u^j$}
               }
          $j \gets j+1$
          }
     }
\Return $\gamma$
\caption{Explicit-element($q$)}\label{explicit}
\end{algorithm}
}

\end{document}